OpenStudy (anonymous):

Simplify quantity 4 times x minus 20 all over 4 times x. Answer quantity x minus 5 all over x quantity x minus 5 all over 4 times x negative 20 over 1 quantity x minus 5 all over 4

OpenStudy (anonymous):

(4x - 20)/(4x) = (x - 5)/x just divide top and bottom by 4

OpenStudy (anonymous):

( 4x - 20) / 4x = 4 ( x - 5) / 4x = (x -5)/ x

OpenStudy (anonymous):

quantity x minus 5 all over x